### v1

@@ 7 Conclusion
The sequence form’s holistic property enables the development of a sequence-form characterization for normal-form refinements of Nash equilibria, and its compactness contributes to computational efficiency. Inspired by this, we have developed a sequence-form characterization of normal-form perfect equilibria for extensive-form games with perfect recall. Guided by this theoretical foundation, we have proposed three distinct sequence-form differential path-following methods for computing normal-form perfect equilibria and rigorously proved their convergence. These methods are underpinned by the construction of artificial games, where the first method incorporates logarithmic-barrier terms into the payoff functions, while the last two methods are derived by extending Harsanyi’s linear and logarithmic tracing procedure to the sequence form, respectively. All three methods provide flexibility in choosing the starting point within a specified range. Through both theoretical analysis and numerical experiments, we have demonstrated the existence of smooth paths leading to normal-form perfect equilibria. To compare the performance of these methods, we have designed two distinct types of random games for comparative experiments. The experimental results further substantiate the effectiveness and efficiency of our methods. Future work could investigate the computation of other normal-form refinements of Nash equilibrium in $n$-player games, such as normal-form proper equilibrium.
[…]
Acknowledgments This work was partially supported by GRF: CityU 11306821 of Hong Kong SAR Government.
@@
Applying backward induction to the first group of equations in the system (10), we derive the following equations for $i\in N$, $j\in M_{i}$, and $a\in A(I^{j}_{i})$,
$\begin{array}[]{l}-\nu^{i}_{I^{j}_{i}}+\sum\limits_{w^{i}\in W^{i},w^{i}_{I^{j}_{i}}a\subseteq w^{i}}((1-c(t))g^{i}(w^{i},\gamma^{*-i}(t))+\lambda^{i}(w^{i}))\prod\limits_{a_{q}\in w^{i}\backslash w^{i}_{I^{j}_{i}}a,a_{q}\in A(I^{j_{q}}_{i})}(1-\theta(t))\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=0,\end{array}$ (27)
@@
$\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=\frac{\gamma^{i}(t;w^{i}_{I^{j_{q}}_{i}}a_{q})-\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j_{q}}_{i}}a_{q})}{(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j_{q}}_{i}})+\theta(t)\gamma^{0i}(w^{i}_{I^{j_{q}}_{i}})-\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j_{q}}_{i}})}>0.$
@@
It can be seen that $\sum_{a_{q}\in A(I^{j_{q}}_{i})}\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=1$. Specifically, for $(j,a)\in D_{i}$, the equations (27) follow directly from the first group of equations in the system (10). When $(j,a)\notin D_{i}$, we assume that the equations (27) hold for all $j_{q}\in M_{i}(w^{i}_{I^{j}_{i}}a)$ and $a_{q}\in A(I^{j_{q}}_{i})$. By multiplying both sides of the equations (27) by $\beta^{i}_{I^{j_{q}}_{i}}(a_{q})$ and summing over $a_{q}\in A(I^{j_{q}}_{i})$, we obtain the expression for $\nu^{i}_{I^{j_{q}}_{i}}$. Finally, by substituting $\zeta^{i}_{I^{j}_{i}}(a)$ with the recursive outcomes, the resulting equation (27) for $(j,a)\notin D_{i}$ is derived. For further analysis, we multiply $\beta^{i}_{I^{j}_{i}}(a)$ on both sides of the equation (27) and sum over $a\in A(I^{j}_{i})$, yielding
@@
$\begin{array}[]{l}-\nu^{i}_{I^{j}_{i}}+\sum\limits_{a\in A(I^{j}_{i})}\sum\limits_{w^{i}\in W^{i},w^{i}_{I^{j}_{i}}a\subseteq w^{i}}((1-c(t))g^{i}(w^{i},\gamma^{*-i}(t))\\ \hskip 128.0374pt+\lambda^{i}(w^{i}))\beta^{i}_{I^{j}_{i}}(a)\prod\limits_{a_{q}\in w^{i}\backslash w^{i}_{I^{j}_{i}}a,a_{q}\in A(I^{j_{q}}_{i})}(1-\theta(t))\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=0.\end{array}$ (28)
@@
Let $L^{i}_{0}=\min_{h\in Z}u^{i}(h)$, $U^{i}_{0}=\max_{h\in Z}u^{i}(h)$ and $Y^{i}_{0}=\max_{j\in M_{i},a\in A(I^{j}_{i})}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)/(\gamma^{0i}(w^{i}_{I^{j}_{i}})-\eta^{0i}(w^{i}_{I^{j}_{i}}))$ for $i\in N$. The equations (28) indicate that $\nu^{i}_{I^{j}_{i}}\geq-|L^{i}_{0}|$ for any $i\in N,j\in M_{i}$. Then we proceed to analyze the upper bound of $(\lambda^{*},\nu^{*})$ under two distinct cases.
@@
Case 1 ($\frac{3}{2}\leq t\leq 2$): In this case, the inequality $\frac{1}{3}\leq\theta(t)\leq 1$ holds. From the equations (28) and the third group of equations in the system (10), it follows that $\nu^{i}_{I^{j}_{i}}\leq|U^{i}_{0}|+3|W^{i}|Y^{i}_{0}$ for all $i\in N$ and $j\in M_{i}$. This result further implies, based on the first group of equations in the system (10), that $\lambda^{i}(w^{i}_{I^{j}_{i}}a)\leq|W^{i}||U^{i}_{0}|+3|W^{i}|Y^{i}_{0}+|L^{i}_{0}|+|M_{i}(w^{i}_{I^{j}_{i}}a)||L^{i}_{0}|$ for all $i\in N$, $j\in M_{i}$, and $a\in A(I^{j}_{i})$.
@@
Case 2 ($0<t\leq\frac{3}{2}$): In this case, the inequality $0\leq\theta(t)\leq\frac{1}{3}$ holds. Consider $i\in N$ and $j\in M_{i}$ such that $w^{i}_{I^{j}_{i}}=\emptyset$. Since $\lambda^{i}(w^{i})\beta^{i}_{I^{j}_{i}}(a)\prod_{a_{q}\in w^{i}\backslash w^{i}_{I^{j}_{i}}a,a_{q}\in A(I^{j_{q}}_{i})}(1-\theta(t))\beta^{i}_{I^{j_{q}}_{i}}(a_{q})\leq|W^{i}|Y^{i}_{0}$, it can be drawn from the equations (28) that $\nu^{i}_{I^{j}_{i}}\leq|U^{i}_{0}|+|W^{i}|Y^{i}_{0}\triangleq V^{j}_{i}$. Furthermore, $(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)\leq V^{j}_{i}+|L^{i}_{0}|$ holds based on the first group of equations in the system (10). As $\theta(t)\leq\frac{1}{3}$, it follows that $\nu^{i}_{I^{j_{q}}_{i}}\leq\frac{3}{2}(V^{j}_{i}+|L^{i}_{0}|)+(|M_{i}(w^{i}_{I^{j}_{i}}a)|-1)|L^{i}_{0}|\triangleq V^{j_{q}}_{i}$ for $a\in A(I^{j}_{i})$ and $j_{q}\in M_{i}(w^{i}_{I^{j}_{i}}a)$. Proceeding by forward induction and noting the game’s finiteness, $\nu^{i}_{I^{j}_{i}}$ is bounded above by $V^{i}_{0}=\max_{j_{q}\in M_{i}}V^{j_{q}}_{i}$ for any $i\in N,j\in M_{i}$. Finally, the first group of equations in the system (10) implies that $\lambda^{i}(w^{i}_{I^{j}_{i}}a)\leq V^{i}_{0}+|L^{i}_{0}|+|M_{i}(w^{i}_{I^{j}_{i}}a)||L^{i}_{0}|$ for $i\in N,j\in M_{i}$ and $a\in A(I^{j}_{i})$.
@@
$B_{1}=\left(\begin{array}[]{cccc}{b^{1}_{1}}^{\top}&&&\\ &{b^{2}_{1}}^{\top}&&\\ &&\ddots&\\ &&&{b^{m_{n}}_{n}}^{\top}\end{array}\right)\in\mathbb{R}^{m_{0}\times n_{0}}\text{ with }b^{j}_{i}=(\frac{d}{dx^{i}(w^{i}_{I^{j}_{i}}a)}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a):a\in A(I^{j}_{i}))^{\top}\in\mathbb{R}^{|A(I^{j}_{i})|}.$
The matrix $B_{2}\in\mathbb{R}^{m_{0}\times n_{0}}$ assigns to each element, where both row and column indices correspond to $w^{i}_{I^{j}_{i}}\neq\emptyset,i\in N,j\in M_{i}$, the value $(\theta(t)-1)d\gamma^{i}(x,t;w^{i}_{I^{j}_{i}})/dx^{i}(w^{i}_{I^{j}_{i}})$, with remaining entries zero. The vector $C\in\mathbb{R}^{m_{0}}$ represents the partial derivative of the last $m_{0}$ components of $p(x,t,\nu;\alpha)$ with respect to $t$. Since both $I^{n_{0}\times n_{0}}$ and $B_{1}+B_{2}$ are of full-row rank, the Jacobian matrix $Dp(x,t,\nu;\alpha)$ has full-row rank on $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}\times\mathbb{R}^{n_{0}}$.
@@
$\begin{array}[]{l}\lambda^{i}(x,2;w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(x,2;w^{i}_{I^{j}_{i}}a)-\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,i\in N,j\in M_{i}.\\ \end{array}$
@@
and $F=\text{diag}(d\lambda^{i}(x,2;w^{i}_{I^{j}_{i}}a)/dx^{i}(w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$. By applying row and column operations to $Dp_{\alpha}(x,2,\nu)$, we obtain

### v3

@@ 7 Conclusion
The sequence form’s holistic property enables the development of a sequence-form characterization for normal-form refinements of Nash equilibria, and its compactness contributes to computational efficiency. Inspired by this, we have developed a sequence-form characterization of normal-form perfect equilibria for extensive-form games with perfect recall. Guided by this theoretical foundation, we have proposed three distinct sequence-form differential path-following methods for computing normal-form perfect equilibria and rigorously proved their convergence. These methods are underpinned by the construction of artificial games, where the first method incorporates logarithmic-barrier terms into the payoff functions, while the last two methods are derived by extending Harsanyi’s linear and logarithmic tracing procedure to the sequence form, respectively. Each of the three methods allows for choosing the starting point within a specified range. The existence of smooth paths converging to normal-form perfect equilibria has been established by means of theoretical analysis and numerical experiments. To evaluate the performance of these methods, we have constructed two distinct types of random games for comparative experiments. The experimental results further substantiate the effectiveness and efficiency of our methods. Future work could investigate the computation of other normal-form refinements of Nash equilibrium in $n$-player games, such as normal-form proper equilibrium.
Acknowledgments
This work was partially supported by GRF: CityU 11306821 of Hong Kong SAR Government.
@@ Acknowledgments
By employing backward induction on the first set of equations in the system (10), we obtain the following equations for $i\in N$, $j\in M_{i}$, and $a\in A(I^{j}_{i})$,
$\begin{array}[]{l}-\nu^{i}_{I^{j}_{i}}+\sum\limits_{\varpi^{i}\in W^{i},\varpi^{i}_{I^{j}_{i}}a\subseteq\varpi^{i}}((1-c(t))g^{i}(\varpi^{i},\gamma^{*-i}(t))\\ \hskip 71.13188pt+\lambda^{i}(\varpi^{i}))\prod\limits_{a_{q}\in\varpi^{i}\backslash\varpi^{i}_{I^{j}_{i}}a,a_{q}\in A(I^{j_{q}}_{i})}(1-\theta(t))\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=0,\end{array}$ (26)
@@
$\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=\frac{\gamma^{i}(t;\varpi^{i}_{I^{j_{q}}_{i}}a_{q})-\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j_{q}}_{i}}a_{q})}{(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j_{q}}_{i}})+\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j_{q}}_{i}})-\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j_{q}}_{i}})}>0.$
@@
It can be seen that $\sum_{a_{q}\in A(I^{j_{q}}_{i})}\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=1$. Specifically, for $(j,a)\in D_{i}$, the equations (26) follow directly from the first set of equations in the system (10). When $(j,a)\notin D_{i}$, we assume that the equations (26) hold for all $j_{q}\in M_{i}(\varpi^{i}_{I^{j}_{i}}a)$ and $a_{q}\in A(I^{j_{q}}_{i})$. By multiplying both sides of the equations (26) by $\beta^{i}_{I^{j_{q}}_{i}}(a_{q})$ and summing over $a_{q}\in A(I^{j_{q}}_{i})$, we obtain the expression for $\nu^{i}_{I^{j_{q}}_{i}}$. Finally, by substituting $\zeta^{i}_{I^{j}_{i}}(a)$ with the recursive outcomes, the resulting equation (26) for $(j,a)\notin D_{i}$ is derived. For further analysis, we multiply $\beta^{i}_{I^{j}_{i}}(a)$ on both sides of the equation (26) and sum over $a\in A(I^{j}_{i})$, yielding
@@
$\begin{array}[]{l}-\nu^{i}_{I^{j}_{i}}+\sum\limits_{a\in A(I^{j}_{i})}\sum\limits_{\varpi^{i}\in W^{i},\varpi^{i}_{I^{j}_{i}}a\subseteq\varpi^{i}}((1-c(t))g^{i}(\varpi^{i},\gamma^{*-i}(t))\\ \hskip 71.13188pt+\lambda^{i}(\varpi^{i}))\beta^{i}_{I^{j}_{i}}(a)\prod\limits_{a_{q}\in\varpi^{i}\backslash\varpi^{i}_{I^{j}_{i}}a,a_{q}\in A(I^{j_{q}}_{i})}(1-\theta(t))\beta^{i}_{I^{j_{q}}_{i}}(a_{q})=0.\end{array}$ (27)
@@
Let $L^{i}_{0}=\min_{h\in Z}u^{i}(h)$, $U^{i}_{0}=\max_{h\in Z}u^{i}(h)$ and $Y^{i}_{0}=\max_{j\in M_{i},a\in A(I^{j}_{i})}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)/(\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})-\eta^{0i}(\varpi^{i}_{I^{j}_{i}}))$ for $i\in N$. The equations (27) indicate that $\nu^{i}_{I^{j}_{i}}\geq-|L^{i}_{0}|$ for any $i\in N,j\in M_{i}$. Then we proceed to analyze the upper bound of $(\lambda^{*},\nu^{*})$ under two distinct cases.
@@
Case 1 ($\frac{3}{2}\leq t\leq 2$): In this case, the inequality $\frac{1}{3}\leq\theta(t)\leq 1$ holds. From the equations (27) and the third set of equations in the system (10), it follows that $\nu^{i}_{I^{j}_{i}}\leq|U^{i}_{0}|+3|W^{i}|Y^{i}_{0}$ for all $i\in N$ and $j\in M_{i}$. This result further implies, based on the first set of equations in the system (10), that $\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)\leq|W^{i}||U^{i}_{0}|+3|W^{i}|Y^{i}_{0}+|L^{i}_{0}|+|M_{i}(\varpi^{i}_{I^{j}_{i}}a)||L^{i}_{0}|$ for all $i\in N$, $j\in M_{i}$, and $a\in A(I^{j}_{i})$.
@@
Case 2 ($0<t\leq\frac{3}{2}$): In this case, the inequality $0\leq\theta(t)\leq\frac{1}{3}$ holds. Consider $i\in N$ and $j\in M_{i}$ such that $\varpi^{i}_{I^{j}_{i}}=\emptyset$. Since $\lambda^{i}(\varpi^{i})\beta^{i}_{I^{j}_{i}}(a)\prod_{a_{q}\in\varpi^{i}\backslash\varpi^{i}_{I^{j}_{i}}a,a_{q}\in A(I^{j_{q}}_{i})}(1-\theta(t))\beta^{i}_{I^{j_{q}}_{i}}(a_{q})\leq|W^{i}|Y^{i}_{0}$, it can be drawn from the equations (27) that $\nu^{i}_{I^{j}_{i}}\leq|U^{i}_{0}|+|W^{i}|Y^{i}_{0}\triangleq V^{j}_{i}$. Furthermore, $(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)\leq V^{j}_{i}+|L^{i}_{0}|$ holds based on the first set of equations in the system (10). As $\theta(t)\leq\frac{1}{3}$, it follows that $\nu^{i}_{I^{j_{q}}_{i}}\leq\frac{3}{2}(V^{j}_{i}+|L^{i}_{0}|)+(|M_{i}(\varpi^{i}_{I^{j}_{i}}a)|-1)|L^{i}_{0}|\triangleq V^{j_{q}}_{i}$ for $a\in A(I^{j}_{i})$ and $j_{q}\in M_{i}(\varpi^{i}_{I^{j}_{i}}a)$. Proceeding by forward induction and noting the game’s finiteness, $\nu^{i}_{I^{j}_{i}}$ is bounded above by $V^{i}_{0}=\max_{j_{q}\in M_{i}}V^{j_{q}}_{i}$ for any $i\in N,j\in M_{i}$. Finally, the first set of equations in the system (10) implies that $\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)\leq V^{i}_{0}+|L^{i}_{0}|+|M_{i}(\varpi^{i}_{I^{j}_{i}}a)||L^{i}_{0}|$ for $i\in N,j\in M_{i}$ and $a\in A(I^{j}_{i})$.
@@
$B_{1}=\left(\begin{array}[]{cccc}{b^{1}_{1}}^{\top}&&&\\ &{b^{2}_{1}}^{\top}&&\\ &&\ddots&\\ &&&{b^{m_{n}}_{n}}^{\top}\end{array}\right)\in\mathbb{R}^{m_{0}\times n_{0}}$
with
$b^{j}_{i}=(\frac{d}{dx^{i}(\varpi^{i}_{I^{j}_{i}}a)}\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):a\in A(I^{j}_{i}))^{\top}\in\mathbb{R}^{|A(I^{j}_{i})|}.$
The matrix $B_{2}\in\mathbb{R}^{m_{0}\times n_{0}}$ assigns to each element, where both row and column indices correspond to $\varpi^{i}_{I^{j}_{i}}\neq\emptyset,i\in N,j\in M_{i}$, the value $(\theta(t)-1)d\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}})/dx^{i}(\varpi^{i}_{I^{j}_{i}})$, with remaining entries zero. The vector $C\in\mathbb{R}^{m_{0}}$ represents the partial derivative of the last $m_{0}$ components of $p(x,t,\nu;\alpha)$ with respect to $t$. Since both $I^{n_{0}\times n_{0}}$ and $B_{1}+B_{2}$ are of full-row rank, the Jacobian matrix $Dp(x,t,\nu;\alpha)$ has full-row rank on $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}\times\mathbb{R}^{n_{0}}$.
@@
[…]
@@
and $F=\text{diag}(d\lambda^{i}(x,2;\varpi^{i}_{I^{j}_{i}}a)/dx^{i}(\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$. By applying row and column operations to $Dp_{\alpha}(x,2,\nu)$, we obtain